# Neutrinos from explosive transients at the dawn of multi-messenger astronomy

### v1

@@ 1 Introduction
Rapid progress has occurred in neutrino astronomy in the past decade, with the discovery of astrophysical neutrinos with energy above $\mathcal{O}(10)$ TeV from the IceCube Neutrino Observatory, as well as the growing number of alerts due to likely joint detections of electromagnetic and neutrino signals from blazars, active galaxies (such as NGC 1068), and our own Galaxy 2; 3. In addition, the first detection of gravitational waves from the merger of two neutron stars in 2017 was followed up by electromagnetic signals across wavebands coming from the associated gamma-ray burst and the kilonova (i.e., the bright electromagnetic transient powered by the radioactive decay of isotopes heavier than iron synthesized in the merger) 4; 5. Neutrinos were not detected in coincidence with the gravitational wave event GW 170817 6; 7, in agreement with theoretical expectations. However, neutrinos are essential characters in the physics of neutron-star merger remnants, indirectly affecting the multi-messenger emission 8. The modeling of neutrino physics and weak interactions is swiftly advancing in light of our renewed understanding of the impact that such physics has on the remnants of neutron-star mergers as well as in the explosive death of massive stars, and nucleosynthesis 9; 10; 11; 12; 13.
@@
The growing number of multi-messenger detections calls for an advanced modeling of the source and related particle production, as in most cases inconsistencies arise between naïve theoretical models and observations 14; 15; 16. In addition, the advent of time-domain astronomy has led to the discovery of new classes of astrophysical transients with puzzling features, such as fast blue optical transients 17; 18, which could be neutrino emitters 19. These sources are characterized by a lightcurve with a rise time faster than the typical supernova one, high peak luminosity, and their photon spectra appear blue. The understanding of the physics governing these objects is crucial to understand how massive stars end their lives and whether these sources could act as additional reservoirs of chemical elements.
@@
In the light of these exciting major theoretical developments and the advent of large-scale observational facilities, this review outlines the state-of-the-art in neutrino astrophysics within a multi-messenger framework, it is arranged as follows. In Sec. 2, the theory underlying the collapse of a massive star, the physics of neutron-star mergers, as well as recent developments on the role of neutrinos in these sources are discussed. The features of the detectable thermal [$\mathcal{O}(10)$ MeV] neutrino signal are also explored, together with what we could potentially learn from such detections. Section 3 focuses on the non-thermal, high-energy [$10^{3}$$10^{9}$ GeV] neutrino emission from sources stemming from collapsing massive stars. The source regions where efficient particle acceleration could take place are highlighted, jointly with how the detection of high-energy neutrinos could be instrumental to discriminate the processes powering the source. Multi-messenger detection strategies are outlined. Finally, future challenges and opportunities are discussed in Sec. 4.
@@ 2.1.1 The core-collapse supernova mechanism
According to the neutrino-driven mechanism 20; 21, about one percent of the binding energy is deposited by neutrinos (the gravitational binding energy of the collapsing massive star is $\sim 3\times 10^{53}$ ergs, with $99\%$ of this energy being carried by neutrinos) in the region behind the stagnant shock through electron (anti)neutrino captures on free neutrons (protons) until matter is heated up so much as to overcome the ram pressure of the infalling layers. Large scale asymmetries form in the postshock region due to hydrodynamic instabilities (e.g., convective overturns of neutrino-heated gas and the standing accretion shock instability, SASI) 22; 23. Such instabilities support the shock revival, pushing it outwards and enabling the runaway shock expansion responsible for the supernova explosion. Neutrinos also drive their hydrodynamical instability during the stalled-shock phase: the lepton-number emission self-sustained asymmetry (LESA) 24. In the presence of LESA, the flux of electron neutrinos minus antineutrinos develops a large-scale emission asymmetry, with implications for neutron-star kicks, flavor conversion, and the neutron-to-proton ratio in the neutrino-heated outflows.
@@
The extreme speed observed for pulsars 25 hints that neutron stars should have natal kicks due to asymmetric ejection of matter and neutrinos. Hydrodynamic simulations of the core collapse show that hydrodynamic and neutrino kicks (e.g. due to LESA) contribute to the natal kick of the compact remnant 26; 27. The recent electromagnetic detection of the inert (i.e. X-ray dormant) stellar-mass black hole binary VFTS 243 allows for the first direct inference of the neutrino natal kick 28.
@@
According to the magneto-rotational mechanism 29, if the proto-neutron star is fast rotating, its rotational energy can be tapped by magnetic fields. The latter can be amplified by convective dynamo, turbulent flows, and magnetorotational instabilities until the energy required to revive the shock is achieved. In some cases, jets can be launched, driving highly asymmetric stellar explosions. If the jet launching is not successful, the jet may be choked in the dense stellar envelope, with no or dim electromagnetic signatures. Explosions driven by rotation and magnetic fields are expected to lead to superluminous supernovae and certain gamma-ray bursts 30; 31. It remains to be understood up to which level assumptions on the pre-collapse magnetic field profile could affect the post-collapse dynamics of the massive star.
@@
During the stellar collapse, we expect neutrinos of all flavors to be emitted for $\mathcal{O}(10)$ s and with an average energy of $\mathcal{O}(10)$ MeV 32 (cf. Table 1 for a summary of the multi-messenger emission from a core-collapse supernova; an overview of the neutrino absorption, production, and direction-changing processes shaping the neutrino distributions in core-collapse supernovae is provided in 33; 34). Gravitational waves should be emitted approximately in the same timeframe as MeV neutrinos. The neutronization burst of electron neutrinos is expected to be visible for the first tens of milliseconds. Then, during the accretion phase, electron neutrinos and antineutrinos are emitted with roughly similar luminosities, which are larger than the non-electron neutrino ones. As the newly born proto-neutron star cools and deleptonizes, after the explosion, (anti)neutrinos of all flavors are emitted with comparable luminosity. The antineutrino fluence at Earth from a core-collapse supernova in our Galaxy is shown in blue in the left panel of Fig. 1 as a function of the neutrino energy; for simplicity, flavor conversion (see Sec. 2.1.2) is neglected; the fluence at Earth should be a combination of the non-oscillated ones represented in Fig. 1.
@@
Type IIp SN $4.4\times 10^{4}$ 35 $10^{42}$ 36 $10^{53}$ 32 $3.3\times 10^{40}$ 37 $10^{44}$$10^{48}$ 38
@@
Type IIn SN $8\times 10^{3}$ 35 $10^{43}$ 39 $3.3\times 10^{41}$ 37
@@
Type Ib/c SN $3\times 10^{4}$ 40 $10^{43}$ 41 $6.7\times 10^{33}$ 37
@@
NS merger & 101010101700170017001700 45 $10^{41}$$10^{42}$ 46 $1.5\times 10^{53}$ 11 $5\times 10^{47}$ 47 $\lesssim 2.2\times 10^{53}$ 48
@@
Figure 1: Antineutrino fluence from a supernova and a neutron-star merger (left panel) and cumulative flux of neutrinos from all supernovae in our Universe (right panel) . Left: Fluences of $\bar{\nu}_{e}$ (solid lines) and $\bar{\nu}_{x}$ ($=\bar{\nu}_{\mu}$ or $\bar{\nu}_{\tau}$, dashed lines) from a core-collapse supernova (in blue) and a neutron-star merger remnant (in orange, multiplied by $4$ for visual purposes), both located at $10$ kpc from Earth. All curves have been obtained neglecting flavor conversion. The core-collapse supernova model is a $20\leavevmode\nobreak\ M_{\odot}$ progenitor 60 with SFHo nuclear equation of state 61. The merger remnant model is an asymmetric merger of two neutron stars with masses of $1.2$ and $1.5\leavevmode\nobreak\ M_{\odot}$ and SFHo nuclear equation of state 62. For our models, the $\bar{\nu}_{e}$ fluence from neutron-star merger remnants is about one order of magnitude smaller than the supernova one; however, neutron-star mergers are not expected to occur as frequently as supernovae. Right: Diffuse supernova neutrino background in the electron antineutrino channel as a function of the neutrino energy. The band is representative and obtained allowing for uncertanties on the supernova rate and fraction of black hole forming collapses 63. The black error bars represent the upper limits from Super-Kamiokande loaded with Gadolinium 64. The DSNB detection is approaching, with the Super-Kamiokande limits touching the theoretical predictions.
The SN 1987A remains the only core-collapse supernova detected in neutrinos up to date 65. While only approximately two dozen neutrino events were gathered in 1987, these data have been crucial showing broad agreement with theoretical expectations; the observed duration of the neutrino burst provided the first evidence of neutrino diffusion out of a dense core. Recent work has given a fresh look into our understanding of the collected neutrino events, confirming that the theoretical models agree well with the SN 1987A neutrino signal detected during the first seconds 66. However, the neutrino events observed around $10$ s are still challenging to explain, since state-of-the-art supernova simulations including proto-neutron-star convection and nucleon correlations in the opacities of neutrinos predict a shorter cooling phase.
@@ 2.1.2 Flavor conversion of neutrinos
[ couleur=bgblue, arrondi=0, logo= , barre=none, noborder=true]Neutrino quantum kinetics When the density of neutrinos is large, as it is the case in the proximity of the proto-neutron star, neutrinos interact among themselves, in addition to interactions with the matter background. To model this phenomenon, for a time $t$, location $\mathbf{x}$, and momentum $\mathbf{p}$, the neutrino mean field is represented through a $3\times 3$ density matrix in the flavor basis [${\mathcal{\rho}}(t,\mathbf{x},\mathbf{p})$ for neutrinos and ${\bar{\varrho}}(t,\mathbf{x},\mathbf{p})$ for antineutrinos]. The diagonal elements of the density matrix represent the occupation numbers linked to the phase-space densities and the off-diagonal elements track flavor coherence. The equations of motion for neutrinos and antineutrinos are, respectively 67:
@@
Due to the large neutrino density in the proximity of the neutrino decoupling region, the coherent forward scattering of neutrinos on each other can trigger flavor conversion, if a crossing in the angular distributions of the electron neutrino lepton number occurs. The conversion rate depends on the (anti)neutrino number density and has been dubbed “fast,” since this flavor conversion can potentially exist in the limit of vanishing neutrino masses 69; 70; this is different from the “slow” neutrino-neutrino flavor conversion that requires neutrino mass splittings.
Investigations relying on hydrodynamical simulations of the core collapse suggest that crossings in the neutrino lepton number could occur even before neutrino decoupling 71; 72 (note, however, that the flavor solution would require a consistent evolution with the supernova hydrodynamics, and this could affect the formation of crossings). Before neutrinos stream freely, they may also undergo flavor conversion triggered by collisional instabilities, if the collisional rate between neutrinos and antineutrinos should be different enough 73. Potentially, flavor conversion induced by collisional instabilities could interfere with fast flavor conversion, although recent work suggests that, in the presence of crossings in the electron neutrino lepton number, the growth rate of the fast flavor instability is larger than the one due to collisional instabilities 74; 75. Yet, non-forward collisions have a non-negligible impact on flavor conversion, enhancing or suppressing the latter according to the shape of the angular distribution of the lepton number of neutrinos, or operating together with flavor conversion to shape the emerging neutrino spectra 76; 77. MSW conversion is expected to further modify the flavor distribution in the outer layers of the star, beyond the shock radius. A visual summary of the flavor conversion regions in the supernova interiors is displayed in the left panel of Fig. 2. The combination of neutrino-neutrino and MSW flavor conversions is expected to mold the neutrino signal observed at Earth 78, as well as the nucleosynthesis in the neutrino-driven outflows 79.
@@
The numerical solution of the seven-dimensional neutrino quantum kinetic equations coupled to the supernova hydrodynamics is not yet available given the technical and conceptual challenges caused by the large difference in characteristic scales of the terms entering the neutrino equations of motion 80. The collapse of massive stars is currently modeled through three-dimensional hydrodynamical simulations, where neutrino flavor conversion is neglected 81. Such simplification was justified in the past, also because it was a common understanding that flavor conversion could only take place beyond the shock radius, with no implications on the explosion mechanism 82. A recent growing body of work points towards the possibility that flavor conversion triggered by neutrino-neutrino interaction can alter the neutrino distributions of all flavors in the source core, even before neutrinos decouple from matter 83. By altering the flavor ratio within the gain region, flavor conversion could aid or hinder the supernova explosion mechanism according to the region where flavor conversion kicks in 9; 10.

### v2

@@ 1 Introduction
Rapid progress has occurred in neutrino astronomy in the past decade, with the discovery of astrophysical neutrinos with energy above $\mathcal{O}(10)$ TeV from the IceCube Neutrino Observatory 2, as well as the growing number of likely joint detections of electromagnetic and neutrino signals from blazars (e.g. TXS 0506+056) 3, active galaxies (such as NGC 1068) 4, and our own Galaxy 5. In addition, the first detection of gravitational waves from the merger of two neutron stars in 2017 was followed up by electromagnetic signals across wavebands coming from the associated gamma-ray burst and the kilonova (i.e., the bright electromagnetic transient powered by the radioactive decay of isotopes heavier than iron synthesized in the merger) 6; 7. Neutrinos were not detected in coincidence with the gravitational-wave event GW 170817 8; 9, in agreement with theoretical expectations. However, neutrinos are essential characters in the physics of neutron-star merger remnants, indirectly affecting the multi-messenger emission 10. The modeling of neutrino physics and weak interactions is swiftly advancing in light of our renewed understanding of the impact that such physics has on the remnants of neutron-star mergers as well as in the explosive death of massive stars, and nucleosynthesis 11; 12; 13; 14; 15.
@@
The growing number of multi-messenger detections calls for an advanced modeling of the source and related particle production, as in most cases inconsistencies arise between naïve theoretical models and observations 16; 17; 18. In addition, the advent of time-domain astronomy has led to the discovery of new classes of astrophysical transients with puzzling features, such as fast blue optical transients 19; 20, which could be neutrino emitters 21; these sources are characterized by a lightcurve with a rise time faster than the typical supernova one, high peak luminosity, and their photon spectra appear blue. The multi-messenger signals arising from the engulfment of a neutron star or a black hole by a giant star will be useful to shed light on the mechanisms powering subsets of observed transients 22; 23; 24; 25; 26. The understanding of the physics governing these objects is crucial to grasp how massive stars end their lives and whether these sources could act as additional reservoirs of chemical elements.
@@
In light of these exciting major theoretical developments and the advent of large-scale observational facilities, this review outlines the state-of-the-art in neutrino astrophysics within a multi-messenger framework, it is arranged as follows. In Sec. 2, the theory underlying the collapse of a massive star, the physics of neutron-star mergers, as well as recent developments on the role of neutrinos in these sources are discussed. The features of the detectable thermal [$\mathcal{O}(10)$ MeV] neutrino signal are also explored, together with what we could potentially learn from such detections. Section 3 focuses on the non-thermal, high-energy [$10^{3}$$10^{9}$ GeV] neutrino emission from sources stemming from collapsing massive stars. The source regions where efficient particle acceleration could take place are highlighted, jointly with how the detection of high-energy neutrinos could be instrumental to discriminate the processes powering the source. Multi-messenger detection strategies are outlined. Finally, future challenges and opportunities are discussed in Sec. 4.
@@ 2.1.1 The core-collapse supernova mechanism
According to the neutrino-driven mechanism 27; 28, about one percent of the binding energy is deposited by neutrinos (the gravitational binding energy of the collapsing massive star is $\sim 3\times 10^{53}$ ergs, with $99\%$ of this energy being carried by neutrinos) in the region behind the stagnant shock through electron (anti)neutrino captures on free neutrons (protons) until matter is heated up so much as to overcome the ram pressure of the infalling layers. Large scale asymmetries form in the postshock region due to hydrodynamic instabilities (e.g., convective overturns of neutrino-heated gas and the standing accretion shock instability, SASI) 29; 30. Such instabilities support the shock revival, pushing it outwards and enabling the runaway shock expansion responsible for the supernova explosion. Neutrinos also drive their hydrodynamical instability during the stalled-shock phase: the lepton-number emission self-sustained asymmetry (LESA) 31. In the presence of LESA, the flux of electron neutrinos minus antineutrinos develops a large-scale emission asymmetry, with implications for neutron-star kicks, flavor conversion, and the neutron-to-proton ratio in the neutrino-heated outflows.
@@
The extreme speed observed for pulsars 32 hints that neutron stars should have natal kicks due to asymmetric ejection of matter and neutrinos. Hydrodynamic simulations of the core collapse show that hydrodynamic and neutrino kicks (e.g. due to LESA) contribute to the natal kick of the compact remnant 33; 34. The recent electromagnetic detection of the inert (i.e. X-ray dormant) stellar-mass black hole binary VFTS 243 allows for the first direct inference of the (neutrino) natal kick 35.
@@
According to the magneto-rotational mechanism 36, if the proto-neutron star is fast rotating, its rotational energy can be tapped by magnetic fields. The latter can be amplified by convective dynamo, turbulent flows, and magnetorotational instabilities until the energy required to revive the shock is achieved. In some cases, jets can be launched, driving highly asymmetric stellar explosions. If the jet launching is not successful, the jet may be choked in the dense stellar envelope, with no or dim electromagnetic signatures. Explosions driven by rotation and magnetic fields are expected to lead to superluminous supernovae and certain gamma-ray bursts 37; 38. It remains to be understood up to which level assumptions on the pre-collapse magnetic field profile could affect the post-collapse dynamics of the massive star.
@@
During the stellar collapse, we expect neutrinos of all flavors to be emitted for $\mathcal{O}(10)$ s and with an average energy of $\mathcal{O}(10)$ MeV 39 (cf. Table 1 for a summary of the multi-messenger emission from a core-collapse supernova; an overview of the neutrino absorption, production, and direction-changing processes shaping the neutrino distributions in core-collapse supernovae is provided in 40; 41). Gravitational waves should be emitted approximately in the same timeframe as MeV neutrinos. The neutronization burst of electron neutrinos is expected to be visible for the first tens of milliseconds. Then, during the accretion phase, electron neutrinos and antineutrinos are emitted with roughly similar luminosities, which are larger than the non-electron neutrino ones. As the newly born proto-neutron star cools and deleptonizes, after the explosion, (anti)neutrinos of all flavors are emitted with comparable luminosity. The antineutrino fluence at Earth from a core-collapse supernova in our Galaxy is shown in blue in the left panel of Fig. 1 as a function of the neutrino energy; for simplicity, flavor conversion (see Sec. 2.1.2) is neglected; the fluence at Earth should be a combination of the non-oscillated ones represented in Fig. 1.
@@
Type IIp SN $4.4\times 10^{4}$ 42 $10^{42}$ 43 $10^{53}$ 39 $3.3\times 10^{40}$ 44 $10^{44}$$10^{48}$ 45
@@
Type IIn SN $8\times 10^{3}$ 42 $10^{43}$ 46 $3.3\times 10^{41}$ 44
@@
Type Ib/c SN $3\times 10^{4}$ 47 $10^{43}$ 48 $6.7\times 10^{33}$ 44
@@
NS merger & 101010101700170017001700 52 $10^{41}$$10^{42}$ 53 $1.5\times 10^{53}$ 13 $5\times 10^{47}$ 54 $\lesssim 2.2\times 10^{53}$ 55
@@
Figure 1: Antineutrino fluence from a supernova and a neutron-star merger (left panel) and cumulative flux of neutrinos from all supernovae in our Universe (right panel) . Left: Fluences of $\bar{\nu}_{e}$ (solid lines) and $\bar{\nu}_{x}$ ($=\bar{\nu}_{\mu}$ or $\bar{\nu}_{\tau}$, dashed lines) from a core-collapse supernova (in blue) and a neutron-star merger remnant (in orange, multiplied by $4$ for visual purposes), both located at $10$ kpc from Earth. All curves have been obtained neglecting flavor conversion. The core-collapse supernova model is a $20\leavevmode\nobreak\ M_{\odot}$ progenitor 67 with SFHo nuclear equation of state 68. The merger remnant model is an asymmetric merger of two neutron stars with masses of $1.2$ and $1.5\leavevmode\nobreak\ M_{\odot}$ and SFHo nuclear equation of state 69. For our models, the $\bar{\nu}_{e}$ fluence from neutron-star merger remnants is about one order of magnitude smaller than the supernova one; however, neutron-star mergers are not expected to occur as frequently as supernovae. Right: Diffuse supernova neutrino background in the electron antineutrino channel as a function of the neutrino energy. The band is representative and obtained allowing for uncertanties on the supernova rate and fraction of black hole forming collapses 70. The black error bars represent the upper limits from Super-Kamiokande loaded with Gadolinium 71. The detection of the DSNB is expected in the near future, with the Super-Kamiokande limits touching the theoretical predictions.
The SN 1987A remains the only core-collapse supernova detected in neutrinos up to date 72; 73; 74. While only approximately two dozen neutrino events were gathered in 1987, these data have been crucial showing broad agreement with theoretical expectations; the observed duration of the neutrino burst provided the first evidence of neutrino diffusion out of a dense core. Recent work has given a fresh look into our understanding of the collected neutrino events, confirming that the theoretical models agree well with the SN 1987A neutrino signal detected during the first seconds 75. However, the neutrino events observed around $10$ s are still challenging to explain, since state-of-the-art supernova simulations including proto-neutron-star convection and nucleon correlations in the opacities of neutrinos predict a shorter cooling phase.
@@ 2.1.2 Flavor conversion of neutrinos
[ couleur=bgblue, arrondi=0, logo= , barre=none, noborder=true]Neutrino quantum kinetics When the density of neutrinos is large, as it is the case in the proximity of the proto-neutron star, neutrinos interact among themselves, in addition to interactions with the matter background. To model this phenomenon, for a time $t$, location $\mathbf{x}$, and momentum $\mathbf{p}$, the neutrino mean field is represented through a $3\times 3$ density matrix in the flavor basis [${\mathcal{\rho}}(t,\mathbf{x},\mathbf{p})$ for neutrinos and ${\bar{\varrho}}(t,\mathbf{x},\mathbf{p})$ for antineutrinos]. The diagonal elements of the density matrix represent the occupation numbers linked to the phase-space densities and the off-diagonal elements track flavor coherence. The equations of motion for neutrinos and antineutrinos are, respectively 76:
@@
Due to the large neutrino density in the proximity of the neutrino decoupling region, the coherent forward scattering of neutrinos on each other can trigger flavor conversion, if a crossing in the angular distributions of the electron neutrino lepton number occurs. The conversion rate depends on the (anti)neutrino number density and has been dubbed “fast,” since this flavor conversion can potentially exist in the limit of vanishing neutrino masses 78; 79; this is different from the “slow” neutrino-neutrino flavor conversion that requires neutrino mass splittings.
Investigations relying on hydrodynamical simulations of the core collapse suggest that crossings in the neutrino lepton number could occur even before neutrino decoupling 80; 81 (note, however, that the flavor solution would require a consistent evolution with the supernova hydrodynamics, and this could affect the formation of crossings). Before neutrinos stream freely, they may also undergo flavor conversion triggered by collisional instabilities, if the collisional rate between neutrinos and antineutrinos should be different enough 82. Potentially, flavor conversion induced by collisional instabilities could interfere with fast flavor conversion, although recent work suggests that, in the presence of crossings in the electron neutrino lepton number, the growth rate of the fast flavor instability is larger than the one due to collisional instabilities 83; 84. Yet, non-forward collisions have a non-negligible impact on flavor conversion, enhancing or suppressing the latter according to the shape of the angular distribution of the lepton number of neutrinos, or operating together with flavor conversion to shape the emerging neutrino spectra 85; 86. MSW conversion is expected to further modify the flavor distribution in the outer layers of the star, beyond the shock radius. A visual summary of the flavor conversion regions in the supernova interiors is displayed in the left panel of Fig. 2. The combination of neutrino-neutrino and MSW flavor conversions is expected to mold the neutrino signal observed at Earth 87, as well as the nucleosynthesis in the neutrino-driven outflows 88.
@@
The numerical solution of the seven-dimensional neutrino quantum kinetic equations coupled to the supernova hydrodynamics is not yet available given the technical and conceptual challenges caused by the large difference in characteristic scales of the terms entering the neutrino equations of motion 89. The collapse of massive stars is currently modeled through three-dimensional hydrodynamical simulations, where neutrino flavor conversion is neglected 90; 91; 28. Such simplification was justified in the past, also because it was a common understanding that flavor conversion could only take place beyond the shock radius, with no implications on the explosion mechanism 92. A recent growing body of work points towards the possibility that flavor conversion triggered by neutrino-neutrino interaction can alter the neutrino distributions of all flavors in the source core, even before neutrinos decouple from matter 93. By altering the flavor ratio within the gain region, flavor conversion could aid or hinder the supernova explosion mechanism according to the region where flavor conversion kicks in 11; 12.